# Investigating the transverse-momentum- and pseudorapidity-dependent flow vector decorrelation in p–Pb collisions with a Multi-Phase Transport model

### v1

@@ Abstract
The event-by-event fluctuations in the initial energy density of the nuclear collisions lead to the decorrelation of second order flow vector, as known as its transverse-momentum ($p_{\mathrm{T}}$) and pseudorapidity ($\eta$) dependence as observed in high-energy heavy-ion collisions. Existing measurements at the CERN Large Hadron Collider shown that these decorrelations are also observed in small collision systems. In this work, a systematic study of the transverse-momentum- and pseudorapidity-dependent flow vector decorrelation is performed in p–Pb collisions at the 5.02 TeV with A Multi-Phase Transport (AMPT) model using different tunings of the initial conditions, partonic and hadronic interactions. It is found that the string-melting version of the AMPT model provides a reasonable description of the measured flow vector decorrelation as a function of $p_{\mathrm{T}}$ and $\eta$. We demonstrate that the hadronic scatterings do not have significant impact on decorrelation in p–Pb collisions for different centrality selections, while both initial conditions and partonic interactions influence the magnitude of the decorrelations. In addition, we found that the subtraction of the nonflow, especially the long-range jet correlation, is crucial for the accurate extraction of the flow vector decorrelation in small collision systems. The comparison of data and model presented in this paper provide further insights in understanding the fluctuations of the flow vector with $p_{\mathrm{T}}$ and $\eta$ in small collision systems and has referential value for future measurements.
@@ I Introduction
High-energy heavy-ion collisions, such as those at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) [1, 2, 3, 4, 5, 6, 7, 8], offer an unique opportunity to study the nuclear matter under extreme conditions of temperature and density. These collisions create the quark-gluon plasma (QGP) [9, 10], a state of matter where quarks and gluons are no longer confined within hadrons but instead form a hot, dense medium. One of the most important observable for the formation of the QGP is the azimuthal anisotropy of produced particles, which is typically characterized by the Fourier expansion of the hadron yield distribution as a function of azimuthal angle $\varphi$:
@@
Due to the event-by-event fluctuations in the initial energy density of the nuclear collisions [33, 29], the fluctuations of second order flow vector $V_{2}$, as well as its transverse-momentum ($\it{p}_{\mathrm{T}}$) dependence have been shown in hydrodynamical models [34, 35, 36]. These fluctuations could result in the breakdown of factorization of two-particle angular correlations into a product of single-particle flow coefficients in different $\it{p}_{\mathrm{T}}$ intervals [37, 34, 38, 39, 40], which were discovered in Pb–Pb collisions by the ALICE and CMS Collaborations [41, 42, 43, 44]. In addition, the flow vector fluctuations along the pseudorapidity ($\eta$) direction was investigated by hydrodynamic and parton transport models, where the factorization breakdown in $\eta$ was found to be sensitive to event-plane fluctuations at different $\eta$ [45, 46, 47, 48]. The LHC experiments measured the ratio of two-particle Fourier coefficients in rapidity bins $\eta$ and $-\eta$ [42, 49, 50], and the results show that the longitudinal fluctuations lead to a linear decrease of the ratio with $\eta$, i.e. the longitudinal flow decorrelations. These studies provide further constraint on the initial conditions and new insights to the longitudinal evolution of the medium formed in heavy ion collisions.
@@
In recent years, the observation of non-zero $v_{2}$ in p–Pb and pp collisions [51, 52, 53, 54, 55, 56, 57, 58, 59] raised the question of whether hydrodynamic flow also exists in these small collision system. The extracted flow harmonics in p–Pb collisions have been studied in detail as a function of $\it{p}_{\mathrm{T}}$, $\eta$ and event multiplicity [60, 54]. On the other hand, the fluctuations of flow vector as a function of $\it{p}_{\mathrm{T}}$ and $\eta$ in small collision systems are measured by the LHC experiments [42, 41, 50, 57]. The effect of factorization breakdown was observed in p–Pb collisions and much smaller than that in Pb–Pb collisions. The measurements of longitudinal flow decorrelations were even extended to the smaller pp collisions by the ATLAS Collaboration [57], and the results reveal significant sensitivity to the nonflow correlations. Since the origin of collective flow in small collision systems remains unclear [61], there are significant uncertainties in the study of $\it{p}_{\mathrm{T}}$ and $\eta$-dependent flow vector decorrelations in these systems. The study based on the event-by-event hydrodynamic calculations demonstrated the important role of longitudinal decorrelations in small collision systems, and predicted rapidity-asymmetric flow decorrelations across different colliding energies [62, 63, 64, 65, 66]. On the other hand, the transpot model has been successfully applied to describe collective phenomena in small collision systems, including charged-particle elliptic flow ($v_{2}$), triangular flow ($v_{3}$[67], and flavor-dependent $\it{p}_{\mathrm{T}}$-differential flow [68, 69, 70, 71, 72, 73]. Building on these successful applications, we present the first systematic study of the $\it{p}_{\mathrm{T}}$- and $\eta$-dependent flow vector decorrelation in p–Pb collisions using the AMPT (A Multi-Phase Transport) model.
@@
The paper is organized as follows. In section II, a brief introduction about the AMPT and its various configurations is given. Section III introduces the observables to characterize the flow vector decorrelation as a funcyion of the $\it{p}_{\mathrm{T}}$ and $\eta$, respectively. The two-particle correlation method and the advanced nonflow subtraction strategy are also presented. The results and related discussions are in section IV. Finally, a summary of this work is presented in section V.
@@ II THE MODEL
In order to separate the impact from different physical processes on the flow vector decorrelation, we varied the key parameters of the model in the calculations of the flow vector. The effect of the partonic phase is investigated by varying the partonic scattering cross section $\sigma$ from 0 mb and 0.5 mb to 3 mb. A smaller cross section corresponds to higher shear viscosity to entropy density ratio ($\eta/s$) in viscous hydrodynamics in the AMPT model, and 0 mb of the cross section represents the exclusion of the parton scattering effect. Following the previous studies about the elliptic flow in p–Pb collisions [70, 71, 72], the default settings of Lund string parameters are $a=0.3$ and $b=0.15$. In this work, we further studied the initial conditions by varying $a=0.5$ and $b=0.9$, corresponding to a smaller string tension, when the partonic scattering cross section $\sigma$ is set to 0.5 mb. In addition, the cutoff time $t_{\mathrm{max}}$ is set to 0.6 fm/$\it{c}$ to turn off the hadronic rescatterings, while the resonance decays are still included. We also analyzed the hadrons obtained right after the quark coalescence in the AMPT evolution (i.e. $t_{\mathrm{max}}=0$ fm/$\it{c}$), where both hadronic rescatterings and resonance decays are disabled. All configurations are summarized in the table 1.
Lund a Lund b Partonic cross section (mb) $t_{\mathrm{max}}$ fm/$\it{c}$
@@ III OBSERVABLES AND METHOD
The flow vector decorrelation is studied using the two-particle correlation method, which was widely applied in previous $v_{2}$ measurements in p–Pb collisions [52, 79]. To simplify, the azimuthal correlation between two emission particles can be represented by $N^{\mathrm{pair}}$ particle pairs as a function of the relative azimuthal angle $\Delta\varphi=\varphi_{a}-\varphi_{b}$ between particles $a$ and $b$ and expanded in the Fourier series as follows:
@@
$C(\Delta\varphi)=\frac{\mathrm{d}N^{\mathrm{pair}}}{d\Delta\varphi}\propto 1+2\sum_{n=1}^{\infty}V_{n\Delta}(\it{p}_{\rm T}^{a},\it{p}_{\rm T}^{b})\mathrm{cos}[n(\Delta\varphi)],$ (3)
@@
where $V_{n\Delta}$ refers to the two-particle $n$-th order harmonic. The tradition approach to calculate the $\it{p}_{\mathrm{T}}$-differential flow of particle $a$ (i.e. the particle of interest, POIs) is to firstly determine the flow coefficient of reference particles (RPs, labeld as ”ref”) over a wide kinematic range, called reference flow, and then the flow coefficient from $\it{p}_{\mathrm{T}}^{a}$ interval can be expressed as:
$v_{n}\{2\}(p_{\mathrm{T}}^{a})$ $=\frac{V_{n\Delta}(p_{\mathrm{T}}^{a},\mathrm{ref})}{\sqrt{V_{n\Delta}(\mathrm{ref},\mathrm{ref})}}$ (4)
$=\frac{\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}^{\mathrm{ref}}\mathrm{cos}[n(\mathrm{\Psi}_{n}(p_{\mathrm{T}}^{a})-\mathrm{\Psi}_{\mathrm{n}})]\rangle}{\sqrt{\langle{v_{n}^{\mathrm{ref}}}^{2}\rangle}},$
where $V_{n\Delta}(p_{\mathrm{T}}^{a},\mathrm{ref})$ and $V_{n\Delta}(\mathrm{ref},\mathrm{ref})$ are obtained by construting the two particle correlation function $C(\Delta\varphi)$ as Eq. 3 between the POIs and RPs, and the RPs, respectively. The $\mathrm{\Psi}_{n}(p_{\mathrm{T}}^{a})$ represents the $p_{\mathrm{T}}$-differential symmetry plane angles at a specific $p_{\mathrm{T}}$ range, which might fluctuate around the reference symmetry plane angles $\mathrm{\Psi}_{n}$. The effect of the difference between $\mathrm{\Psi}_{n}(p_{\mathrm{T}}^{a})$ and $\mathrm{\Psi}_{n}$ is quantified by the cosine term $\mathrm{cos}[n(\Psi_{n}(p_{\mathrm{T}}^{a})-\Psi_{n})]$, which is called $\it{p}_{\mathrm{T}}$-dependent flow angle fluctuations. On the other hand, the fluctuations of the flow magnitude are observed when the factorization hypothesis is broken [43, 44], i.e. the $\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}^{\mathrm{ref}}\rangle$ cannot be factorised into the product of $\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle}$ and $\sqrt{\langle{v_{n}^{\mathrm{ref}}}^{2}\rangle}$. Another $\it{p}_{\mathrm{T}}$ differential flow observable, denoted as $v_{n}[2](p_{\mathrm{T}}^{2})$, was proposed in  [80], which is not affected by fluctuations in the flow angle or flow magnitude:
$v_{n}[2](p_{\mathrm{T}}^{a})=\sqrt{V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{a})}=\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle}.$ (5)
@@
The ratio of $v_{n}\{2\}(p_{\mathrm{T}}^{a})$ and $v_{n}[2](p_{\mathrm{T}}^{a})$ is proposed to probe the $p_{\mathrm{T}}$-dependent flow vector fluctuations:
$\frac{v_{n}\{2\}(p_{\mathrm{T}}^{a})}{v_{n}[2](p_{\mathrm{T}}^{a})}=\frac{\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}^{\mathrm{ref}}\mathrm{cos}[n(\mathrm{\Psi}_{n}(p_{\mathrm{T}}^{a})-\mathrm{\Psi}_{\mathrm{n}})]\rangle}{\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle}\sqrt{\langle{v_{n}^{\mathrm{ref}}}^{2}\rangle}}.$ (6)
@@
One can see that if the $p_{\mathrm{T}}$-dependent flow vector fluctuations present, the ratio value is smaller than unity. Similarly, another observable to probe the $p_{\mathrm{T}}$-differential factorization ratio $r_{n}$ was proposed in [81, 80],
$r_{n}$ $=\frac{V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{b})}{\sqrt{V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{a})V_{n\Delta}(p_{\mathrm{T}}^{b},p_{\mathrm{T}}^{b})}}$ (7)
$=\frac{\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}(p_{\mathrm{T}}^{b})\mathrm{cos}[n(\Psi_{n}(p_{\mathrm{T}}^{a})-\Psi_{n}(p_{\mathrm{T}}^{b}))]\rangle}{\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle\langle v_{n}(p_{\mathrm{T}}^{b})^{2}\rangle}}$
@@
The $r_{n}<$ 1 indicates the breaking of factorization, suggesting the presence of $p_{\mathrm{T}}$-dependent flow vector fluctuations.
@@
Figure 1: (Color online) The $v_{2}\{2\}$ and $v_{2}[2]$ as a function of $p_{\mathrm{T}}$ and their ratio obtained from the AMPT model in 0–20%, 20–40% and 40–60% p–Pb collisions at 5.02 TeV, in comparison of the ALICE measurements [41]. The results with the nonflow subtraction obtained from the AMPT model are also shown.
Similarly, the $\eta$-dependent flow harmonics $V_{n\Delta}(\eta^{a},\eta^{b})$ are calculated by constructing two-particle correlations between pseudorapidity bins $\eta^{a}$ and $\eta^{b}$, following the same method as Eq. 3 but replacing $p_{T}^{\mathrm{a}}$ and $p_{T}^{\mathrm{b}}$ with $\eta^{a}$ and $\eta^{b}$. In order to investigate the longitudinal fluctuations of the flow vector, the decorrelation ratio $r_{n}(\eta^{a},\eta^{b})$ was proposed in Ref. [42], which is defined as:
$r_{n}(\eta^{a},\eta^{b})$ $=\frac{V_{n\Delta}(-\eta^{a},\eta^{b})}{V_{n\Delta}(\eta^{a},\eta^{b})}$ (8)
$=\frac{\langle v_{n}(-\eta^{a})v_{n}(\eta^{b})\mathrm{cos}[n(\Psi_{n}(-\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle}{\langle v_{n}(\eta^{a})v_{n}(\eta^{b})\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle},$
where the $\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]$ represents the impact of the variation in symmetry plane angles between the $\eta^{a}$ and $\eta^{b}$ bins, resulting from the $\eta$-dependent flow angle fluctuations. In addition, the flow magnitude decorrelation introduces additional $\eta$-dependent fluctuations, lead to the break of factorization of flow coefficient products: $\langle v_{n}(-\eta^{a})v_{n}(\eta^{b})\rangle\neq\sqrt{\langle v_{n}^{2}(-\eta^{a})\rangle}\sqrt{\langle v_{n}^{2}(\eta^{b})\rangle}$. For symmetric collision systems like Pb–Pb collisions, the $v_{n}$ obtained from symmetric positive $[v_{n}(\eta_{a})]$ and negative $[v_{n}(-\eta_{a})]$ $\eta$ regions are identical after averaging over all events. Therefore, the $r_{n}(\eta^{a},\eta^{b})$ is expected to converge to unity if neither flow magniture nor flow angle fluctuate along the longitudinal direction. If either of these effects, or both, are present, the $r_{n}(\eta^{a},\eta^{b})$ will be less than unity [50]. However, in an asymmetric collision system, such as p–Pb collisions, $v_{n}(\eta_{a})$ and $v_{n}(-\eta_{a})$ are generally not identical. As a result, it is not possible to isolate the $\eta$-dependent effects of symmetry-plane fluctuations in Eq. 8. The CMS collaboration proposed a method [42] to take the product of $r_{n}(\eta^{a},\eta^{b})$ and $r_{n}(-\eta^{a},-\eta^{b})$:
$\sqrt{r_{n}(\eta^{a},\eta^{b})r_{n}(-\eta^{a},-\eta^{b})}$ (9)
$=\sqrt{\frac{\langle v_{n}(-\eta^{a})v_{n}(\eta^{b})\rangle\langle v_{n}(\eta^{a})v_{n}(-\eta^{b})\rangle}{\langle v_{n}(\eta^{a})v_{n}(\eta^{b})\rangle\langle v_{n}(-\eta^{a})v_{n}(-\eta^{b})\rangle}}\cdot$
$\sqrt{\frac{\langle\mathrm{cos}[n(\Psi_{n}(-\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle}{\langle\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle}\frac{\langle\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(-\eta^{b}))]\rangle}{\langle\mathrm{cos}[n(\Psi_{n}(-\eta^{a})-\Psi_{n}(-\eta^{b}))]\rangle}}$
@@
In this case, the $v_{n}$ terms can be canceled in if there is no flow vector magnitude fluctuations, then the $\eta$-dependent flow vector decorrelation in p–Pb collisions can also be studied.
@@
Figure 2: (Color online) The $p_{\mathrm{T}}$-dependent $v_{2}\{2\}$/$v_{2}[2]$ ratio in 0–20%, 20–40%, 40–60% centrality classes from the AMPT model with different configurations. ALICE data points [41] are shown as black open circles for comparison.
(a) The AMPT calculations are performed with different cross setions $\sigma$ of parton scatterings.
@@
Figure 3: (Color online) The factorization ratio $r_{2}$ as a function of $p^{a}_{\mathrm{T}}$ obtained from the AMPT calculations, in different centrality and $p^{b}_{\mathrm{T}}$ ranges, is compared to the ALICE data points [41].
(a) The AMPT calculations are performed with different cutoff time $\tau$ and Lund parameters $a$ and $b$.
Figure 4: (Color online) The factorization ratio $r_{2}$ as a function of $p^{a}_{\mathrm{T}}$ obtained from the AMPT calculations, in different centrality and $p^{b}_{\mathrm{T}}$ ranges, is compared to the ALICE data points [41].
@@
In small collision systems, the main contributions of the nonflow is from the jets, including both the short-range and long-range (also called ”recoil”) jet correlations [82]. The former can be effectively removed by introducing a large rapidity gap between the trigger and associated particles during the construction of the correlations, and the latter can be suppressed by many methods in the measurements based on the different assumptions. In this work, the template fit method developed by the ATLAS collaboration [60] is applied, which has been proven effective for nonflow subtraction in our previous studies [70] in the AMPT. For the given $\it{p}_{\mathrm{T}}$ and $\eta$ integral, the correlation function distribution obtained in high-multiplicity events is assumed to result from the superposition of the distribution obtained in low-multiplicity events scaled up by a multiplicative factor $F$ and a constant modulated by $\mathrm{cos}(n\Delta\varphi)$ for $n>1$, as shown in
$C(\Delta\varphi)=FC^{\mathrm{LM}}(\Delta\varphi)+G(1+2\sum_{n=1}^{3}V_{n\Delta}\mathrm{cos}(n\Delta\varphi)),$ (10)

### v2

# Probing transverse-momentum- and pseudorapidity-dependent flow vector decorrelation in p–Pb collisions at the CERN Large Hadron Collider

@@ Abstract
The event-by-event fluctuations in the initial energy density of the nuclear collisions lead to the decorrelation of second order flow vector, as well as its transverse-momentum ($p_{\mathrm{T}}$) and pseudorapidity ($\eta$) dependence as observed in high-energy heavy-ion collisions. Recent measurements at the CERN Large Hadron Collider shown that these decorrelations are also observed in small collision systems. In this work, a systematic study of the transverse-momentum- and pseudorapidity-dependent flow vector decorrelation is performed in p–Pb collisions at the 5.02 TeV with A Multi-Phase Transport (AMPT) model using different tunings of the initial conditions, partonic and hadronic interactions. It is found that the string-melting version of the AMPT model provides a reasonable description of the measured flow vector decorrelation as a function of $p_{\mathrm{T}}$ and $\eta$. We demonstrate that the hadronic scatterings do not have significant impact on decorrelation in p–Pb collisions for different centrality selections, while both initial conditions and partonic interactions influence the magnitude of the decorrelations. In addition, the nonflow, especially the long-range jet correlation, plays an important role in generating the decorrelation in small collision systems. The comparison of data and model presented in this paper provide further insights in understanding the fluctuations of the flow vector with $p_{\mathrm{T}}$ and $\eta$ in small collision systems and has referential value for future measurements.
@@ I Introduction
High-energy heavy-ion collisions, such as those at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) [32, 17, 19, 41], offer an unique opportunity to study the nuclear matter under extreme conditions of temperature and density. These collisions create the quark-gluon plasma (QGP) [45, 46], a state of matter where quarks and gluons are no longer confined within hadrons but instead form a hot, dense medium. One of the most important observable for the formation of the QGP is the azimuthal anisotropy of produced particles, which is typically characterized by the Fourier expansion of the hadron yield distribution as a function of azimuthal angle $\varphi$:
@@
Due to the event-by-event fluctuations in the initial energy density of the nuclear collisions [52, 35], the fluctuations of second order flow vector $V_{2}$, as well as its transverse-momentum ($\it{p}_{\mathrm{T}}$) dependence have been shown in hydrodynamical models [24, 23]. These fluctuations could result in the breakdown of factorization of two-particle angular correlations into a product of single-particle flow coefficients in different $\it{p}_{\mathrm{T}}$ intervals [42, 24, 30, 56], which were discovered in Pb–Pb collisions by the ALICE and CMS Collaborations [11, 36]. In addition, the flow vector fluctuations along the pseudorapidity ($\eta$) direction was investigated by hydrodynamic and parton transport models, where the factorization breakdown in $\eta$ was found to be sensitive to event-plane fluctuations at different $\eta$ [43, 22, 53, 21]. The CMS Collaboration measured the ratio of two-particle Fourier coefficients in rapidity bins $\eta$ and $\eta$ [36], and the results show that the longitudinal fluctuations lead to a linear decrease of the ratio with $\eta$, i.e. the longitudinal flow decorrelations. These studies provide further constraint on the initial conditions and new insights to the longitudinal evolution of the medium formed in heavy ion collisions.
@@
In recent years, the observation of non-zero $v_{2}$ in p–Pb and pp collisions raised the question of whether hydrodynamic flow also exist in these small collision systems [29, 9, 5]. The extracted flow harmonics in p–Pb collisions have been studied in detail as a function of $\it{p}_{\mathrm{T}}$, $\eta$ and event multiplicity [25, 49, 50, 55]. On the other hand, the fluctuations of flow vector as a function of $\it{p}_{\mathrm{T}}$ and $\eta$ are measured by the ALICE and CMS Collaborations [36, 11]. The effect of factorization breakdown was observed in p–Pb collisions and much smaller than that in Pb–Pb collisions. The measurements of longitudinal flow decorrelations were even extended to the smaller pp collisions by the ATLAS Collaboration [1], and the results reveal significant sensitivity to the nonflow correlations. However, a systematic theoretical study for the $\it{p}_{\mathrm{T}}$- and $\eta$-dependent flow vector decorrelations in small collision systems is still missing. The effects from the initial and final state effects, as well as the nonflow contribution are also needed to be further investigated in the p–Pb collisions. Therefore, we present the first transport model study on this topic.
@@
In this paper, the $\it{p}_{\mathrm{T}}$- and $\eta$-dependent flow vector decorrelation in p–Pb collisions are systematically studied with A Multi-Phase Transport (AMPT) model. The model with different tunings is applied to disentangle the impact of various physical processes, such as initial conditions, parton transport, nonflow contribution, and hadronic rescatterings. In section II, a brief introduction about the AMPT and its various configurations is given. Section III introduces the observables to characterize the flow vector decorrelation as a funcyion of the $\it{p}_{\mathrm{T}}$ and $\eta$, respectively. The two-particle correlation method and the advanced nonflow subtraction strategy are also presented. The results and related discussions are in section IV. Finally, a summary of this work is presented in section V.
@@ II THE MODEL
[…]
@@ III OBSERVABLES AND METHOD
The flow vector decorrelation in p–Pb collisions is investigated by using the two-particle correlation method. To simplify, the azimuthal correlation between two emission particles can be represented by $N^{\mathrm{pair}}$ particle pairs as a function of the relative angle $\Delta\varphi=\varphi_{a}-\varphi_{b}$ between particles a and b and expanded in the Fourier series as follows:
@@
$aligned C(\Delta\varphi)=\frac{\mathrm{d}N^{\mathrm{pair}}}{d\Delta\varphi}\propto 1+2\sum_{n=1}^{\infty}V_{n\Delta}(\it{p}_{\rm T}^{a},\it{p}_{\rm T}^{b})\mathrm{cos}[n(\Delta\varphi)],$ (3)
where $V_{n\Delta}$ refers to the two-particle $n$-th order harmonic. In a pure hydrodynamic scenario, the connection between the single- and two-particle harmonics can be expressed as:
$aligned\langle V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{b})\rangle=\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}(p_{\mathrm{T}}^{b})\mathrm{cos}[n(\Psi_{n}(p_{\mathrm{T}}^{a})-\Psi_{n}({p_{\mathrm{T}}^{b}}))]\rangle,$ (4)
where the bracket $\langle\cdot\rangle$ represents the average over all events, and the cosine term $\mathrm{cos}[n(\Psi_{n}(p_{\mathrm{T}}^{a})-\Psi_{n}({p_{\mathrm{T}}^{b}}))]$ shows the effects of the difference between the symmetry plane angle at $\it{p}_{\mathrm{T}}^{a}$ and $p_{\mathrm{T}}^{b}$ bin, due to the $\it{p}_{\mathrm{T}}$-dependent flow angle fluctuations.
@@
The tradition approach to calculate the $\it{p}_{\mathrm{T}}$-differential flow of particle $a$ (i.e. the particle of interest, POIs) is to firstly determine the flow coefficient of reference particles (RPs) over a wide kinematic range, called reference flow, and then the flow coefficient from $\it{p}_{\mathrm{T}}^{a}$ interval can be expressed as:
$aligned v_{n}\{2\}(p_{\mathrm{T}}^{a})&=\frac{V_{n\Delta}(p_{\mathrm{T}}^{a},\mathrm{ref})}{\sqrt{V_{n\Delta}(\mathrm{ref},\mathrm{ref})}}\\ &=\frac{\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}^{\mathrm{ref}}\mathrm{cos}[n(\mathrm{\Psi}_{n}(p_{\mathrm{T}}^{a})-\mathrm{\Psi}_{\mathrm{n}})]\rangle}{\sqrt{\langle{v_{n}^{\mathrm{ref}}}^{2}\rangle}},$ (5)
where the $v_{n}(p_{\mathrm{T}}^{a})v_{n}^{\mathrm{ref}}$ cannot be factorised into the product of $\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle}$ and $\sqrt{\langle{v_{n}^{\mathrm{ref}}}^{2}\rangle}$ if there are $p_{\mathrm{T}}$-dependent flow vector fluctuations. Another $\it{p}_{\mathrm{T}}$ differential flow observable, denoted as $v_{n}[2](p_{\mathrm{T}}^{2})$, was proposed in ref. [34], which is not affected by the flow vector fluctuations:
$aligned v_{n}[2](p_{\mathrm{T}}^{a})=\sqrt{V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{a})}=\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle}.$ (6)
@@
The ratio of $v_{n}\{2\}(p_{\mathrm{T}}^{a})$ and $v_{n}[2](p_{\mathrm{T}}^{a})$ is proposed to probe the $p_{\mathrm{T}}$-dependent flow vector decorrelation
[…]
@@
One can see that if the $p_{\mathrm{T}}$-dependent flow vector fluctuations present, the ratio value is smaller than unity. Similarly, another observable to probe the $p_{\mathrm{T}}$-differential factorization ratio $r_{n}$ was proposed [31, 34],
$aligned r_{n}&=\frac{V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{b})}{\sqrt{V_{n\Delta}(p_{\mathrm{T}}^{a},p_{\mathrm{T}}^{a})V_{n\Delta}(p_{\mathrm{T}}^{b},p_{\mathrm{T}}^{b})}}\\ &=\frac{\langle v_{n}(p_{\mathrm{T}}^{a})v_{n}(p_{\mathrm{T}}^{b})\mathrm{cos}[n(\Psi_{n}(p_{\mathrm{T}}^{a})-\Psi_{n}(p_{\mathrm{T}}^{b}))]\rangle}{\sqrt{\langle v_{n}(p_{\mathrm{T}}^{a})^{2}\rangle\langle v_{n}(p_{\mathrm{T}}^{b})^{2}\rangle}}$ (8)
@@
The $r_{n}$ ¡ 1 indicates the breaking of factorization, suggesting the presence of $p_{\mathrm{T}}$-dependent flow vector fluctuations.
@@
Figure 1: (Color online) The $v_{2}\{2\}$ and $v_{2}[2]$ as a function of $p_{\mathrm{T}}$ and their ratio obtained from the AMPT model in 0–20%, 20–40% and 40–60% p–Pb collisions at 5.02 TeV, in comparison of the ALICE measurements. The results with the nonflow subtraction obtained from the AMPT model are also shown.
Similarly, the method to calculate the $\eta$-dependent flow is to construct the two-particle correlation between two different pseudo-rapidity bins, and the two-particle harmonics can be expressed as a formalism similar to Eq. 4 by replacing $p_{T}^{a}$ and $p_{T}^{b}$ by $\eta^{a}$ and $\eta^{b}$:
$aligned\langle V_{n\Delta}(\eta^{a},\eta^{b})\rangle=\langle v_{n}(\eta^{a})v_{n}(\eta^{b})\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle,$ (9)
where the $\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]$ represents the impact of the variation in symmetry plane angles between the $\eta^{a}$ and $\eta^{b}$ bins, resulting from the $\eta$-dependent flow vector fluctuations. Then $\eta$-dependent factorization can be studied by the observable $r_{n}(\eta^{a},\eta^{b})$ proposed in Ref[], which is defined as
$aligned r_{n}(\eta^{a},\eta^{b})&=\frac{V_{n\Delta}(-\eta^{a},\eta^{b})}{V_{n\Delta}(\eta^{a},\eta^{b})}\\ &=\frac{v_{n}(-\eta^{a})v_{n}(\eta^{b})\langle\mathrm{cos}[n(\Psi_{n}(-\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle}{v_{n}(\eta^{a})v_{n}(\eta^{b})\langle\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle}.$ (10)
For symmetric collision systems like Pb–Pb collisions, the $v_{n}$ obtained from symmetric positive $v_{n}(\eta_{a})]$ and negative $v_{n}(-\eta_{a})]$ $\eta$ regions are identical after averaging over all events. Therefore, Eq. 10 can be approximated by
$aligned r_{n}(\eta^{a},\eta^{b})\approx\frac{\langle\mathrm{cos}[n(\Psi_{n}(-\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle}{\langle\mathrm{cos}[n(\Psi_{n}(\eta^{a})-\Psi_{n}(\eta^{b}))]\rangle},$ (11)
where the $r_{n}(\eta^{a},\eta^{b})$ is less than unity if the factorization break down (i.e. the decorrelation presents). However, in an asymmetric collision system, such as p–Pb collisions, $v_{n}(\eta_{a})$ and $v_{n}(-\eta_{a})$ are generally not identical. As a result, it is not possible to isolate the $\eta$-dependent effects of symmetry-plane fluctuations in Eq. 10. The CMS collaboration proposed a method to take the product of $r_{n}(\eta^{a},\eta^{b})$ and $r_{n}(-\eta^{a},-\eta^{b})$ to remove the $v_{n}$ terms:
[…]
@@
In this case, the $\eta$-dependent flow vector decorrelation in p–Pb collisions can also be studied.
@@
Figure 2: (Color online) The $p_{\mathrm{T}}$-dependent $v_{2}\{2\}$$v_{2}[2]$ ratio in 0–20%, 20–40%, 40–60% centrality classes from the AMPT model with different configurations. ALICE data points are shown as black open circles for comparison.
@@
Figure 3: (Color online) The factorization ratio $r_{2}$ as a function of $p^{a}_{\mathrm{T}}$, in different centrality and $p^{b}_{\mathrm{T}}$ ranges, obtained from the AMPT model with various configurations. ALICE data points are shown as black open circles for comparison.
@@
In small collision systems, the main contributions of the nonflow is from the jets, including both the short-range and long-range (also called ”recoil”) jet correlations. The former can be effectively removed by introducing a large rapidity gap between the trigger and associated particles during the construction of the correlations, and the latter can be suppressed by many methods in the measurements based on the different assumptions. In this work, the template fit method developed by the ATLAS collaboration is applied, which has been proven effective for nonflow subtraction in our previous studies in the AMPT. For the given $\it{p}_{\mathrm{T}}$ and $\eta$ integral, the correlation function distribution obtained in high-multiplicity events is assumed to result from the superposition of the distribution obtained in low-multiplicity events scaled up by a multiplicative factor $F$ and a constant modulated by $\mathrm{cos}(n\Delta\varphi)$ for $n>1$, as shown in
$aligned C(\Delta\varphi)=FC^{\mathrm{LM}}(\Delta\varphi)+G(1+2\sum_{n=1}^{3}V_{n\Delta}\mathrm{cos}(n\Delta\varphi)),$ (13)